4 3 2 2 = 119 ; 6 5 5 4 = 40 ; 8 4 3 7 = ? Solution please

Grade:10

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

Please send the pic of the source of the question

Please consider this
4*4 + 3-2+2= 19

6*6 +5-5+5= 40

TTherefore 8*8 +4-3+7= 72

Regards